10.1 A study reports that the mean birth weight of 81 full-term infants is 6.1 pounds. The study also reports SE-0.22 pounds. 95 a) What is the margin of error of this estimate for 95% confidence? b) What is the 95% confidence interval for ? c) What does it mean when we say that we have 95% confidence in the interval? d) Recall that SEX-in. Rearrange this equation to solve for σ. What was the standard deviation of birth weights in this population?
Answer #1

10.1 A study reports that the mean birth weight of 81 full-term infants is 6.1 pounds the study also reports ''SE = 0.22 pounds.

(a)What is the marginal error of this estimate for 95% confidence.   

Answer:- so that Margin of Error at 95% Interval = 1.960*0.22 = 0.43119 Pounds

(b) what is the 95% confidence interval for meu.

Answer:-

Interval for mean 95%

(6.1-0.43119, 6.1+0.43119)

= (5.67, 6.53)

c) We are 95% Confident that the Mean birth weight of full term newborn children will lies in this interim

(d) σ 0.22 * V81 = 1.98

please do rate or liked.............